OpenStudy (anonymous):

how do i eliminate the parameter that has the same terms in both x and y..i.e two cosθ's

OpenStudy (anonymous):

\[x=1+\cosθ \]

OpenStudy (anonymous):

how do i eliminate the parameter, would it be come an inverse?

OpenStudy (anonymous):

\[r \cos \theta = x\] \[r \sin \theta = y\] \[r ^{2}=a ^{2}+b ^{2}\] Use these substitutions.

OpenStudy (anonymous):

r^2 = x^2 + y^2 I mean, sorry. (its the Pythagorean theorem, thats why I mixed up the letters!)